# MEAL: A Benchmark for Continual Multi-Agent Reinforcement Learning

### v1

@@ 5 Experiments
Figure 3: Average Normalize Score over the course of training on a sequence of 20 randomly generated Level 1 tasks. Shaded regions indicate 95% confidence intervals across 5 seeds.
@@ 5.1 Setup
The agent is trained on each task $\mathcal{T}_{i}$ for $\Delta=10^{7}$ environment steps on-policy with the dense reward setting, repeated over five seeds. During training, we evaluate the policy after every 100 updates by running 10 evaluation episodes on all previously seen tasks. To enable fair comparisons across tasks with different layouts, sizes, and object configurations, we normalize raw returns by dividing the total number of soups delivered by a programmatically estimated upper bound specific to each layout (see Appendix A.1). This bound estimates the maximum number of soups a single agent could deliver in an episode by executing the most efficient cook–deliver cycle. A normalized score of 1 indicates the agent matches this single-agent optimum, while values above 1 indicate successful cooperation between agents that exceeds solo efficiency. We leverage JAX to reduce the wall-clock time for training on a single environment to around 5 minutes. All experiments are conducted on a dedicated compute node with a 72-core 3.2 GHz AMD EPYC 7F72 CPU and a single NVIDIA A100 GPU. We adopt many of JaxMARL’s default settings for our network configuration, IPPO setup, and training processes. For exact hyperparameters please refer to Appendix B.2.
5.2 Baselines
We evaluate several continual learning methods. Fine-Tuning (FT) is a naive baseline where the policy is trained sequentially across tasks without any mechanism to prevent forgetting. L2-Regularization [16] adds a penalty on parameter changes to encourage stability. EWC [16] is a regularization method that penalizes changes to important parameters, with importance measured using the Fisher Information Matrix. Online EWC is a variant that maintains a running estimate of parameter importance, making it more suitable for longer sequences. MAS [3] computes importance based on how parameters influence the policy’s output, rather than gradients. As the MARL baseline, we opt for IPPO [9]. It is a natural choice as it can be seamlessly integrated with all continual learning methods. It has been shown to outperform other MARL approaches on both SMAC [9] and Overcooked [25], making it a strong candidate for evaluating CMARL in a fully observable setting.
@@ 5.3 Baseline Comparison
Figure 4: Normalized evaluation scores of each task in a Level 1 20-task sequence during training. Results are averaged over 5 seeds with the shaded areas depicting 95% confidence intervals.
Table 2: $\mathcal{A}$ denotes final average performance, $\mathcal{F}$ quantifies forgetting, and $\mathcal{P}$ reflects plasticity. EWC and L2 retain knowledge best, with EWC achieving the highest overall performance. FT shows the highest plasticity but forgets almost everything.
Method $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{P}\!\uparrow$
FT 0.048±$\pm 0.00$ 0.899±$\pm 0.01$ 1.163±$\pm 0.03$
Online EWC 0.769±$\pm 0.09$ 0.150±$\pm 0.09$ 1.136±$\pm 0.03$
EWC 0.839±$\pm 0.03$ $\pm 0.03$ 1.062±$\pm 0.01$
MAS 0.281±$\pm 0.07$ 0.501±$\pm 0.06$ 0.914±$\pm 0.07$
L2 0.753±$\pm 0.02$ 0.031±$\pm 0.00$ 0.869±$\pm 0.08$
Figure 3 compares the performance of several continual learning methods combined with IPPO over a 20-task sequence, and Table 2 reports summary metrics over 5 seeds. Fine-Tuning (FT) shows the highest plasticity, but fails catastrophically at retention: once a task is left behind, its performance rapidly collapses. EWC and L2 retain knowledge nearly perfectly, with EWC also achieving the highest average performance. Online EWC matches their plasticity but retains less. MAS ends up in the middle: moderate forgetting and low plasticity, resulting in significantly lower final performance.
5.4 Forgetting
Figure 4 illustrates the extent of forgetting across tasks for Fine-Tuning (FT), EWC, and MAS. FT suffers from clear catastrophic forgetting: once the agent transitions to a new task, performance on the previous task collapses immediately. In contrast, EWC shows near-perfect retention across the sequence, maintaining stable performance on all tasks. MAS falls in between: it neither retains past knowledge effectively nor reaches the same per-task performance as FT or EWC, suggesting it struggles with both plasticity and stability in this setting.
@@ 5.5 Ablation Study
Figure 5: Ablation results for EWC, MAS, and L2 on Level 1 10-task sequences. The error bars indicate the 95% confidence interval across 5 seeds. Removing the multi-head architecture leads to severe performance degradation across all methods. Including the task ID and critic regularization have a negligible effect. Layer normalization has little impact on EWC or MAS, but significantly improves L2. A CNN encoder yields worse results than an MLP variant.
To determine which components are crucial for CMARL on MEAL, we ablate five components in our default IPPO learning setup: multi-head architectures, task identity inputs, critic regularization, layer normalization, and replacing the MLP with a CNN encoder. The results in Figure 5 reveal that multi-head outputs are most critical for MEAL task sequences. Removing them consistently devastates performance across all methods, likely due to uncontrolled interference between tasks in the shared output head. In contrast, not providing the model with the one-hot encoded task ID vector has a negligible effect. Prior continual RL studies [31, 30] report that it is beneficial to only regularize the actor and let the critic adapt freely. In our experiments, however, we find that this has little effect. Layer normalization shows method-specific sensitivity: while it makes little difference for EWC and MAS, it more than doubles the performance of L2 regularization. This is likely because L2 penalizes absolute weight magnitudes, and layer norm helps stabilize activations across tasks, mitigating harmful scale drift. Finally, swapping to a CNN encoder substantially hurts performance for all methods. Given the small layouts in Level 1 tasks (6×6 to 7×7), CNNs struggle to extract meaningful features and add unnecessary parameter overhead, making simple MLPs the better fit in this setting.
@@ 5.6 Network Plasticity
Table 3: Averaged plasticity metrics for FT.
Repeats AUC-loss $\downarrow$ FPR $\uparrow$ RAUC $\uparrow$
3 0.154 1.228 0.905
10 0.215 0.896 0.846
@@
Figure 6: Plasticity degradation in MEAL under fine-tuning (FT). We measure capacity loss (left), final-performance ratio (middle), and raw-AUC ratio (right) over a sequence of 10 tasks. Each line shows the mean across 5 seeds; error bars denote ±$\pm 1$ s.e. Increasing the repetition count from 3 to 10 amplifies degradation across all metrics, indicating reduced plasticity as learning progresses.
@@
A well-documented pitfall in continual RL is the gradual loss of plasticity, an agent’s ability to fit new data after many tasks [1, 10]. To test whether MEAL exhibits the same pathology, we repeat a Level 1 10-task sequence multiple times and compare performance between repetitions. We evaluate the fine-tuning (FT) baseline, as it ranks highest in plasticity according to Table 2. We track three standard metrics, each normalized to the initial repetition: (i) AUC-loss $\downarrow$ captures capacity drop, (ii) Final-performance ratio (FPR) $\uparrow$ compares the final plateau to the first repetition, and (iii) Raw-AUC ratio (RAUC) $\uparrow$ measures total reward accumulated during the repeat. For formal definitions of the metrics and training curves, see Appendix C.2. Figure 6 shows that repeating tasks 10 times leads to a consistently greater loss of plasticity across all metrics, compared to 3 repetitions, in most tasks. Table 3 summarises the sequence-level means: AUC-loss rises $\approx 40\%$, while FPR drops below 1, and RAUC falls $6.5\%$, confirming that MEAL triggers the same degradation seen in single-agent continual RL.
@@ 6 Conclusion
We introduced MEAL, the first benchmark for continual multi-agent reinforcement learning. By leveraging JAX for efficient GPU-accelerated training and introducing a diverse set of handcrafted and procedurally generated Overcooked environments, MEAL enables the study of long-horizon continual learning in cooperative settings. Our evaluation of six continual learning methods combined with the IPPO algorithm reveals that existing CL techniques struggle to retain cooperative behaviors while maintaining adaptability to new tasks. Regularization-based methods mitigate forgetting but sacrifice plasticity, while parameter-isolation methods fail to scale with longer task sequences. These findings highlight the need for new approaches that can handle the dual challenges of cooperation and non-stationarity in CMARL. We hope MEAL serves as a foundation for advancing this underexplored but important research direction.
@@ 7 Limitations
While MEAL provides a scalable and diverse testbed for continual multi-agent reinforcement learning, several limitations remain. First, MEAL is restricted to fully observable, two-agent environments with discrete action spaces, limiting its applicability to partially observable or competitive multi-agent settings. Second, while layout diversity is high, the domain itself is narrow. Overcooked dynamics do not capture the full complexity of real-world multi-agent interactions involving language, negotiation, or long-horizon planning. Third, our benchmark only evaluates task-incremental learning by changing layouts. Future work could extend MEAL to other continual learning settings. Finally, we only consider continual learning in settings where the environment layout changes across tasks, but not the partner agent or environment dynamics.
@@ Appendix A Implementation Details
A.1 Maximum Soup Delivery Estimator
@@
Every pick-up or drop is assumed to take a constant $c_{\text{act}}=2$ steps (turn $+$ interact). With $n_{\text{int}}=3\!\times\!2+1+1+1=9$ interactions per cycle, the overhead is $c_{\text{over}}=n_{\text{int}}\,c_{\text{act}}=18$.
@@
The bound assumes a single agent acting optimally. It ignores multi–agent collaboration and therefore underestimates throughput in layouts where multiple agents can parallelize the workflow. Listing A.1 contains the exact implementation.
{listing}
[!t] Heuristic upper bound (estimate_max_soup). {minted}python # overcooked_upper_bound.py (excerpt) DELIVERY_REWARD = 20 COOK_TIME = 20 ACTION_OVERHEAD = 2 INTERACTIONS_PER_CYCLE = 3 * 2 + 1 + 1 + 1 OVERHEAD_PER_CYCLE = INTERACTIONS_PER_CYCLE * ACTION_OVERHEAD
def estimate_cycle_time(layout, n_agents=2): … move_cost = 3 * d_onion + d_plate + 1 + d_goal + 3 return move_cost + COOK_TIME + OVERHEAD_PER_CYCLE
def estimate_max_soup(layout, episode_len, n_agents=2): cyc = estimate_cycle_time(layout, n_agents) soups = episode_len // cyc return int(soups * DELIVERY_REWARD)

### v2

@@ 5 Experiments
Figure 3: Top: The average normalized score evaluation curves on Level 1 tasks show a notable performance gap across baselines. Middle: Per-task evaluation scores for EWC on Level 1 indicate near-perfect retention. Bottom: EWC manages to outperform the standard IPPO baseline by transferring knowledge forward on most Level 2 tasks, while under-performing in others. The green area between the curves indicates positive forward transfer, while red represents the negative counterpart.
@@ 5.1 Setup
The agent is trained on each task $\mathcal{T}_{i}$ for $\Delta=10^{7}$ environment steps on-policy with the dense reward setting, repeated over five seeds. In our experiments, we adopt the task-incremental continual learning paradigm, in which the task identity is known during both training and evaluation. During training, we evaluate the policy after every 100 updates by running 10 evaluation episodes on all tasks in the sequence. The results are displayed with 95% confidence intervals. We leverage JAX to reduce the wall-clock time for training on a single task to around 5 minutes. Out experiments are conducted on a dedicated compute node with a 72-core 3.2 GHz AMD EPYC 7F72 CPU and a single NVIDIA A100 GPU. We adopt many of JaxMARL’s default settings for our network configuration, IPPO setup, and training processes. For exact hyperparameters please refer to Appendix B.2.
@@ 5.2 Baseline Comparison
We evaluate popular CL methods. Fine-Tuning (FT) is a naive baseline where the policy is trained sequentially across tasks without any mechanism to prevent forgetting. L2-Regularization Kirkpatrick et al. (2017) adds a penalty on parameter changes to promote stability. EWC Kirkpatrick et al. (2017) penalizes changes to important parameters, with importance measured using the Fisher Information Matrix. Online EWC is a variant that maintains a running estimate of parameter importance. MAS Aljundi et al. (2018) computes importance based on how parameters influence the policy’s output, rather than gradients. AGEM Chaudhry et al. (2018) is a replay-based method that projects the current gradient update to avoid interference with past tasks, using a memory buffer of stored experiences. As the default MARL algorithm, we opt for IPPO De Witt et al. (2020). It is a natural choice as it can be seamlessly integrated with all model-free CL methods. Moreover, it has been shown to outperform other MARL approaches on SMAC De Witt et al. (2020) and Overcooked Rutherford et al. (2024), making it a strong candidate for evaluating CMARL in MEAL.
Figure 4: Jointly visualizing forward transfer and forgetting results on Level 1 reveals the classic stability–plasticity trade-off in continual learning.
Figure 5: EWC’s performance notably declines as layout complexity increases. Most high-level tasks remain unsolved.
Table 2: Baseline comparison results across three difficulty levels. The confidence intervals are omitted for brevity, see Appendix J for the full results.
Method Level 1 Level 2 Level 3
$\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$ $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$ $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
FT 0.048 0.899 0.201 0.041 0.790 0.065 0.010 0.519 -0.157
EWC 0.839 0.031 0.055 0.604 0.061 -0.086 0.178 0.053 -0.650
Online EWC 0.769 0.150 0.208 0.585 0.211 0.152 0.306 0.168 -0.149
MAS 0.281 0.501 -0.233 0.155 0.423 -0.355 0.034 0.245 -0.542
L2 0.753 0.031 -0.199 0.496 0.106 -0.527 0.127 0.052 -0.827
AGEM 0.204 0.761 0.125 0.117 0.625 -0.083 0.037 0.457 -0.169
Figure 3 (top) compares our baselines on Level 1, and Table 2 reports the exact metrics for all levels. Fine-Tuning (FT) and AGEM show higher forward transfer, but catastrophically fail at retention: once a task is left behind, performance rapidly collapses. EWC and L2 exhibit near-perfect knowledge retention on all levels, with EWC also ranking highest in average score on lower levels. Figure 3 (middle) visualizes EWC’s per-task stability. EWC locks parameters to all past tasks via a cumulative Fisher penalty, which pays off on Levels 1–2, where layouts are small. However, on Level 3, that rigidity bites, as harder layouts demand larger representation shifts, causing EWC to underfit. By contrast, Online EWC uses a decayed Fisher that down-weights older tasks and manages to keep enough plasticity to learn the new layouts (higher $\mathcal{A}$ and $\mathcal{FT}$). See Appendix C for a deeper analysis of this comparison. MAS performs poorly in all metrics, although outperforming FT and AGEM. Notably, the simplistic difficulty level design of MEAL presents notable challenges, as EWC’s score diminishes with increasing difficulty (Figure 5). Figure 5 illustrates the fundamental stability-plasticity trade-off in CL. L2 achieves excellent retention but limited forward transfer, while Fine-Tuning and AGEM demonstrate high plasticity with severe forgetting. EWC and its online version provide a middle ground, balancing both objectives more effectively than other approaches.
@@ 5.3 Ablation Study
Table 3: Comparison of EWC with PPO/IPPO across 1–3 agent task sequences. Two agents yield the best results due to parallelism and cooperative potential. Adding a third agent introduces instability, non-stationarity, and coordination challenges that hurt performance.
Agents Level 1 Level 2 Level 3
$\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$ $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$ $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
1 Agent 0.622 0.046 -0.045 0.343 0.071 -0.458 0.285 0.159 -0.531
2 Agents 0.839 0.031 0.055 0.604 0.061 -0.086 0.178 0.053 -0.650
3 Agents 0.476 0.125 -0.676 0.277 0.132 -0.896 0.117 0.094 -0.978
Figure 6: Ablation results on Level 1 10-task sequences. Removing the multi-head architecture leads to severe performance degradation across all methods. Including the task ID and critic regularization has a negligible effect. Layer normalization has little impact on EWC or MAS, but significantly improves L2. A CNN encoder yields worse results than a simple MLP variant.
To determine which components are crucial for CMARL on MEAL, we ablate five components in our default IPPO learning setup: multi-head architectures, task identity inputs, critic regularization, layer normalization, and replacing the MLP with a CNN encoder. The results in Figure 6 reveal that multi-head outputs are most critical for MEAL task sequences. Removing them consistently devastates performance across all methods, likely due to uncontrolled interference between tasks in the shared output head. In contrast, not providing the model with the one-hot encoded task ID vector has a negligible effect. Prior continual RL studies Wołczyk et al. (2021); Tomilin et al. (2023) report that it is beneficial to only regularize the actor and let the critic adapt freely. In our experiments, however, we find that this has little effect. Layer normalization shows method-specific sensitivity: while it makes little difference for EWC and MAS, it more than doubles the performance of L2 regularization. This is likely because L2 penalizes absolute weight magnitudes, and layer norm helps stabilize activations across tasks, mitigating harmful scale drift. Finally, swapping to a CNN encoder substantially hurts performance for all methods. Given the small layouts in Level 1 tasks (6×6 to 7×7), CNNs struggle to extract meaningful features and add unnecessary parameter overhead, making simple MLPs the better fit in this setting.
5.4 $N$-Agent MEAL
To better analyze the multi-agent dimension of CMARL, we extend MEAL to support an $N$-agent setting, allowing us to systematically study how the number of cooperating agents affects continual learning. We run EWC combined with PPO for the single-agent case, and IPPO for 2 and 3 agents.
A single agent delivers less soup than two because it cannot parallelize tasks: while one agent delivers soup, the other can already load the pot with onions. However, in Level 3, the single-agent PPO outperforms the two-agent IPPO. We observed that, under the CL setting, the two-agent setup solved fewer tasks due to the larger grid size in harder layouts: increasing both the observation space and the number of agents makes the learning problem more complex for IPPO, which struggles to learn a good policy.
Adding a third agent further hurts performance for similar reasons. IPPO trains independent policies while the environment remains a joint MDP, where transitions and rewards depend on the combined actions of all agents. Moving from 2→3 agents expands the joint action space and interaction patterns, amplifies non-stationarity (as two teammates’ policies change simultaneously), and makes credit assignment more difficult (since the reward is shared, IPPO does not know which agent made a good action). Without explicit communication or role allocation, IPPO struggles to learn continually as the team and layout size grow. We explore common pitfalls of agent behavior in Appendix I.
@@ 5.5 Network Plasticity
Table 4: Averaged plasticity metrics for FT on Level 1 task sequences.
Repeats AUCL $\downarrow$ FPR $\uparrow$ RAUC $\uparrow$
1 0.000 1.000 1.000
3 0.154 0.906 0.905
10 0.215 0.866 0.846
Table 5: Effect of reward design on EWC over 20-task Level-1 sequences.
Rewards $\mathcal{A}\!\uparrow$ $\mathcal{F}\!\downarrow$ $\mathcal{FT}\!\uparrow$
Shared 0.90 $\pm$ 0.04 0.01 $\pm$ 0.01 0.20 $\pm$ 0.08
Individual 0.84 $\pm$ 0.08 0.08 $\pm$ 0.07 0.12 $\pm$ 0.06
Sparse 0.19 $\pm$ 0.04 0.02 $\pm$ 0.01 -0.79 $\pm$ 0.10
@@
Figure 7: Plasticity loss in MEAL under fine-tuning (FT). We measure capacity loss (left), final-performance ratio (middle), and raw-AUC ratio (right) over 10-task Level 1 sequences. Increasing the repetition count from 3 to 10 amplifies degradation across all metrics, indicating reduced plasticity as learning progresses. The high error bars stem from the agent learning a task during one repetition but performing poorly during another. This effect is more pronounced at lower repetitions.
@@
A well-documented pitfall in continual RL is the gradual loss of plasticity, an agent’s ability to fit new data after many tasks (Abbas et al., 2023; Dohare et al., 2024). To test whether MEAL exhibits the same pathology, we repeat a Level 1 10-task sequence multiple times and compare performance between repetitions. We evaluate fine-tuning (FT), as it is the simplest and most cost-effective baseline ranking high in plasticity (Figure 5). We track three standard metrics, each normalized to the initial repetition: (i) AUC-loss $\downarrow$ captures capacity drop, (ii) Final-performance ratio (FPR) $\uparrow$ compares the final plateau to the first repetition, and (iii) Raw-AUC ratio (RAUC) $\uparrow$ measures total reward accumulated during the repeat. For formal definitions of the metrics and training curves, see Appendix G.
We observe that all metrics deteriorate with longer training (Table 5 and Figure 7), confirming that loss of plasticity also appears in the multi-agent setting. Notably, the larger drop in performance occurs between 1 and 3 repetitions, suggesting that early degradation is more severe. AUC-loss increases by roughly 40% when going from 3→10 repetitions. Despite our setting spanning over 1B environment steps, well beyond the scale of prior studies (Abbas et al., 2023; Dohare et al., 2024), those works report a much stronger loss of plasticity than observed in MEAL. We hypothesize that this difference stems from our experiments using multiple output heads, which isolate task-specific outputs, reduce gradient interference, and preserve prior policies while allowing the backbone to learn transferable features.
5.6 Reward Design
By default, Overcooked agents receive shared rewards. We compare this with an individual reward mode, where each agent is rewarded solely for its own actions, often leading to greedier behavior and weaker coordination Perolat et al. (2017); Foerster et al. (2017); Hughes et al. (2018). We also evaluate the sparse shared reward setting described in Section 4.1, assessing all three using EWC on 20-task Level 1 sequences.
As shown in Table 5, shared rewards lead to the best performance, although on some tasks, the individual reward setting allows the agents to find a better global solution due to the inherent competitiveness. Agents are motivated to use different onion piles and pots to maximize their own rewards, which often leads to a more efficient solution. However, this competitive drive occasionally prevents them from converging on a stable solution. The sparse reward setting grants rewards only for successful deliveries. Without a targeted exploration mechanism, agents are unlikely to discover a full delivery sequence through random actions even on Level 1 layouts, leading to worse performance.
5.7 Partial Observability
Figure 8: Performance of EWC under full (FO) and partial observability (PO) on 20-task Level 1 sequences. FO yields the best results overall. MAPPO underperforms in the continual learning setting.
Although Overcooked is fully observable by design, we introduce a partially observable variant to better reflect real-world sensing constraints (limited field of view, occlusions). Following popular MARL environments Resnick et al. (2018); Mohanty et al. (2020); Agapiou et al. (2022); Ellis et al. (2023), each agent receives an egocentric, direction-aware observation window with all outside tiles masked. The specification and difficulty scaling of this window are detailed in Appendix D.
In this setting, MAPPO Yu et al. (2022) is known to outperform IPPO, since its centralized critic can 1) more accurately estimate individual contributions to shared rewards under partial observability, and 2) reduce non-stationarity by conditioning value estimates on the joint actions of all agents, leading to more stable and coordinated policy updates. We investigate this by running a 20-task sequence under partial observability (PO) with EWC and compare the results with the fully observable (FO) baseline.
Across all levels, IPPO under full observability (FO) clearly dominates the partial setting (PO), and both outperform MAPPO in our CL regime (Figure 8). The gap between IPPO(FO) and IPPO(PO) stems from full state information simplifying credit assignment and stabilizing value targets. Partial observability thus increases task difficulty. Contrary to expectation, MAPPO underperforms IPPO. A plausible cause is a mismatch between MAPPO’s centralized critic and the CL regime. Conditioning on joint observations and actions drifts substantially across tasks, yielding noisier targets and stronger cross-task interference, while IPPO’s independent critics learn simpler task-local value functions that transfer more stably. These results motivate including PO variants in MEAL to stress coordination under incomplete information for more realistic continual MARL benchmarking.
@@ 6 Conclusion
We introduced MEAL, a scalable benchmark for continual multi-agent RL, built on JAX for efficient GPU training. The on-demand creation of procedurally generated Overcooked layouts enables long-horizon CMARL studies with controlled difficulty, observability, and agent count. We evaluated combinations of popular CL methods and MARL algorithms, revealing that existing techniques struggle to retain cooperative behaviors while maintaining adaptability to new tasks. The $N$-agent extension increases coordination demands and exacerbates interference, yielding a harder, more variable task distribution. Partial observability compounds this difficulty, as centralized critics exhibit stronger cross-task drift and interference. Individual rewards weaken coordination and induce negative transfer. A simple curriculum boosts performance on complex layouts under an equal data budget. Training on long task sequences degrades network plasticity in MARL, while multi-head architectures yield the largest structural gains for performance. Our findings suggest that MEAL exposes the dual challenge of cooperation and non-stationarity in CMARL. We see immediate headroom for methods that (i) are purpose-built for CMARL, jointly handling partner and environment-level non-stationarity, (ii) stabilize credit assignment under partial observability across task sequences, and (iii) drive structured exploration and robust coordination in diverse, long-horizon settings. We hope MEAL serves as a solid foundation for pushing this line of work forward.
@@ 7 Limitations
While MEAL provides a scalable and diverse testbed for CMARL, several limitations remain. First, MEAL is restricted to discrete action spaces, limiting its applicability. Second, while layout diversity is high, the domain itself is narrow. Overcooked dynamics do not capture the full complexity of real-world multi-agent interactions involving language, negotiation, or long-horizon planning. Third, our benchmark only evaluates task-incremental learning by changing layouts. Future work could extend MEAL to other CL settings. Finally, we only consider CL in settings where the environment layout changes across tasks, but not the partner agent nor environment dynamics.
Acknowledgments
This work was conducted with the assistance of the Dutch national e-infrastructure, generously supported by the SURF Cooperative under grant EINF-12816. Additionally, this research was partially funded by the KOIOS project under grant agreement 101103770.
@@ Appendix A Implementation Details
A.1 Maximum Soup Delivery Calculator
@@
Every pick-up or drop is assumed to take a constant $c_{\text{act}}=2$ steps (turn ++ interact). With $n_{\text{int}}=3\!\times\!2+1+1+1=9$ interactions per cycle, the overhead is $c_{\text{over}}=n_{\text{int}}\,c_{\text{act}}=18$.
@@
$R_{\max}(\mathcal{L},H)\;=\;20\,N_{\max}(\mathcal{L},H).$
The bound assumes a single agent acting optimally. It ignores multi–agent collaboration and therefore underestimates throughput in layouts where multiple agents can parallelize the workflow. Listing 9 contains the exact implementation.
Figure 9: Heuristic upper bound (calculate_max_soup).
# overcooked_upper_bound.py (excerpt)
COOK_TIME = 20
ACTION_OVERHEAD = 2
INTERACTIONS_PER_CYCLE = 3 * 2 + 1 + 1 + 1
OVERHEAD_PER_CYCLE = INTERACTIONS_PER_CYCLE * ACTION_OVERHEAD
@@
move_cost = 3 * d_onion + d_plate + 1 + d_goal + 3
return move_cost + COOK_TIME + OVERHEAD_PER_CYCLE
[…]
cyc = calculate_cycle_time(layout, n_agents)
soups = episode_len // cyc
return int(soups)